# SBC collaboration Low-Threshold Response of a Scintillating Xenon Bubble Chamber to Nuclear and Electronic Recoils

### v1

@@ Abstract
A device filled with pure xenon first demonstrated the ability to operate simultaneously as a bubble chamber and scintillation detector in 2017. Initial results from data taken at thermodynamic thresholds down to 4 keV showed sensitivity to 20 keV nuclear recoils with no observable bubble nucleation by -ray interactions. This paper presents results from further operation of the same device at thermodynamic thresholds as low as keV, hardware limited. The bubble chamber has now been shown to have sensitivity to 1 keV nuclear recoils while remaining insensitive to bubble nucleation by -rays. A robust calibration of the chamber’s nuclear recoil nucleation response, as a function of nuclear recoil energy and thermodynamic state, is presented. Stringent upper limits are established for the probability of bubble nucleation by -ray-induced Auger cascades, with a limit of set at keV, the lowest thermodynamic threshold explored.
@@ I Introduction
It was shown in Ref. [12] that pure xenon bubble chambers are sensitive to nuclear recoils, as necessary for a dark matter search, despite being insensitive to -rays. Thus, the scintillating noble liquid bubble chamber is a promising technology for exploring dark matter, in particular (GeV) dark matter where sensitivity to sub-keV nuclear recoils is necessary, as well as for detection of coherent elastic neutrino-nucleus scattering (CENS) [13, 14]. This paper describes further experiments with the same xenon bubble chamber (referred to as “XeBC”) as the one used in Ref. [12], modified for operation at higher degrees of superheat (lower nuclear recoil detection thresholds). A variety of -ray and neutron sources are used to characterize the response of superheated liquid xenon to both electronic and nuclear recoils, as relevant to future dark matter direct detection and CENS experiments.
@@ II Previous Experiments and Motivation
The PICO collaboration has recently published extensive studies of electronic recoil (ER) and nuclear recoil (NR) induced bubble nucleation in C3F8 bubble chambers [15, 16, 17, 18]. The well-established “Seitz” threshold [3, 15] (calculated from the pressure, temperature, and thermophysical properties of the target fluid, and denoted here as ) is used both as a measure of the degree of superheat in these studies, and as an indication of the expected NR detection threshold. In practice, this quantity is seen to underestimate both observed [18, 19, 20, 21, 22, 23] and simulated [24, 25] thresholds for NR-induced bubble nucleation. The level of disagreement varies with both target fluid and threshold, requiring NR nucleation efficiencies to be determined in each new case through dedicated neutron calibrations. The robust approach developed by the PICO collaboration [18] consists of a global fit of simulated to observed bubble nucleation rates with data gathered from multiple neutron calibration sources.
@@
In 2017, XeBC observed for the first time coincident bubble nucleation and scintillation from NRs [12] without the bubble nucleation by photoabsorption seen in xenon-contaminated C3F8. This paper presents results from low-threshold (high-superheat) XeBC data taken in the year following the 2017 publication, as described in Section III, with NR- and ER-analyses in the style of the above PICO analyses presented in Section IV. Section IV.1 uses an approach similar to [18] to constrain the NR nucleation efficiency in xenon at Seitz thresholds between 0.9 and 2.1 keV. Section IV.2 gives upper limits on XeBC’s sensitivity to ER events at Seitz thresholds down to 0.5 keV. No evidence for bubble nucleation by -ray photoabsorption is seen, extending the disparity between xenon and the molecular fluids studied by PICO.
@@ III.2 Chamber Control and Instrumentation
The thermodynamic state (temperature and pressure) of a bubble chamber determines the threshold energy required for bubble nucleation. For XeBC, there are two temperature regions to consider: the warm active region becomes superheated when the chamber pressure drops to the target expanded state, while the jar seal, plumbing, and the bellows that transmits pressure changes from the hydraulic fluid to the target fluid are kept cold so that liquid contacting these surfaces will not boil [16]. The entire chamber assembly is cooled by a cold finger running from an open-mouth dewar of liquid nitrogen to the top of the cold (inactive) region of the bubble chamber. The warm region is then heated with a nichrome wire heater mounted on the outer wall of the copper, aluminum, and sapphire radiation shield that encloses the target volume. This shield is thermally coupled to the outer jar by an indium contact immediately below the target volume. The temperature of the warm region is controlled with feedback from a 100- platinum resistance temperature device (RTD) mounted on the inside of the radiation shield (T2 in Fig. 1). Additional RTDs (T1, T3, T4) monitor temperatures throughout the warm and cold regions.
@@
Pressure is controlled by a hydraulic system filled with Dynalene-MV, a commercial low-temperature / low-viscosity hydrocarbon blend, using a pair of hydraulic accumulators as “reservoirs” for the high (compressed) and low (expanded) pressures. Hydraulic pressure is transmitted to the xenon by an in-line pressure-balanced bellows. A second bellows transmits the pressure back to a warm hydraulic stub, and the pressure of all three volumes (hydraulic system output, xenon, and hydraulic stub) are monitored by absolute pressure transducers with 0.01 bar resolution (PT5, PT4, and PT6, respectively). The redundant pressure readings provide cross-calibrations to compensate for temperature-dependent readings in the cold xenon pressure transducer and for pressure-differentials across the bellows due to the bellows spring constants.
@@
Bubble chamber data is organized by “event”, i.e., one cycle from compressed (insensitive) expanded (superheated, sensitive) trigger compressed. During a run the chamber is set to cycle automatically, repeating the cycle with a 60-s compression period between events. The camera acts as the primary bubble trigger, collecting 0.3-megapixel, 8-bit images at 100 frames per second with synchronized illumination from a ring of 950-nm LEDs mounted around the camera. The frame grabber receiving the camera images broadcasts a trigger when a significant change (typ. 25% full-scale) is seen in a minimum number of pixels (typ. 12 pixels) between consecutive images. In addition to images, data recorded for each event include pressure and temperature readings throughout the event, acoustic waveforms digitized at 2.5 MHz covering 100-ms before and after the bubble trigger, and time-stamped PMT waveforms digitized at 1 GHz for scintillation pulses observed throughout the event. The PMT digitizer is triggered by a NIM discriminator. The PMT trigger is suppressed when camera illumination is on, incurring a 10% dead-time in the scintillation channel, but is otherwise independent from the bubble chamber cycle.
@@ III.3 Low-Threshold Data
The low-threshold data analyzed here were acquired over the course of 5 months from June to October 2017, in three distinct xenon fills. The chamber’s superheated operational temperature ranged from to , with target expansion pressures from 1.7–3.4 bara. Data are reported in terms of the Seitz threshold of the superheated target, calculated following [15] with thermophysical properties for xenon taken from NIST REFPROP [27].
A variety of sources were deployed, including 252Cf as a fission neutron source, 88Y-Be and 207Bi-Be as photoneutron sources, and 88Y, 207Bi, and 57Co as -ray sources. Table 1 summarizes the activities, neutron production rates, placement, and resulting neutron fluxes in the superheated target from these sources as determined by a MCNPX-PoliMi [28] Monte Carlo simulation. The fission source generates neutrons over a broad spectrum with an average energy of 2.3 MeV [29], and is simulated using the built-in PoliMi 252Cf fission spectrum. The photoneutron sources produce nearly mono-energetic low-energy neutrons via 9BeBe reactions in BeO and metallic beryllium parts surrounding the -ray sources. The 88Y-Be source neutrons are peaked at 152 keV, with approximately 0.7% of the neutrons at higher energy (950 keV) from the low-branching-fraction 2734-keV -ray [30]. The 207Bi-Be source produces neutrons peaked at 91 keV. Photoneutron rates and neutron spectra are calculated using MCNPX-PoliMi, using the 9BeBe cross-sections given in [31] and correcting lab-frame neutron energies with the appropriate relativistic kinematic factor (necessary due to MCNPX’s mishandling of momentum [32]).
@@
Table 1 also shows uncertainties in neutron flux from each source, stemming from uncertainties in both source strength and in neutron propagation. The 252Cf source neutron emission rate was calibrated to 6% precision with a well-characterized 3He counter, and was found to be consistent with the high energy (2 MeV) gamma emission rate measured by a NaI crystal. The total uncertainty on neutron flux from the 252Cf source at the chamber is 9% after including geometric uncertainties in neutron propagation, which were estimated by varying selected details in the simulated geometry and by comparing the MCNP-PoliMi simulation with an independently developed Geant4 [33] simulation. The uncertainty in the 88Y-Be source neutron emission rate is 5%, due to a combination of geometrical uncertainties in the 88Y deposition profile in the source capsule and uncertainties on the 9BeBe cross-section, and is consistent with the neutron emission rate measured by the same 3He counter used to characterize the 252Cf source. Neutron propagation uncertainties are significantly higher for 88Y-Be than for 252Cf, giving a total neutron flux uncertainty of 13%. The 207Bi-Be neutron flux uncertainty is larger still at 30%, dominated by uncertainty associated with additional beryllium pieces deployed with that source in an attempt to boost the neutron emission rate. Finally, there is a common-mode 10% uncertainty on the integrated flux of neutrons through the “active” xenon due to uncertainty on the volume of xenon visible to the camera.
@@
Each neutron dataset (source + threshold) is accompanied by a dedicated background measurement taken at the same threshold and in the same xenon fill. In the case of 252Cf this simply involves removing the source. For the 88Y-Be and 207Bi-Be experiments, the Be and BeO components are replaced with aluminum, with the 88Y and 207Bi sources still present to preserve the same -ray flux incident upon the detector but without the photoneutrons.
The 207Bi and 88Y -ray sources without beryllium are also used to measure ER sensitivity, along with the lower energy 57Co source, at Seitz thresholds between 0.50 keV and 1.56 keV. Sources were placed 10 cm from the active xenon for all ER sensitivity measurements.
@@
Reduced data for later analysis include only live-times and bubble counts at each threshold and source configuration (with the exception of 252Cf data, which include the scintillation veto described in the next section). The chamber is considered live only when the measured pressure is within 0.034 bar (0.5 psi) of the target pressure set-point (extended to 0.068 bar for the 0.90 keV 88Y-Be dataset), and any live-time in the the first 20 s of each event is discarded to avoid thermodynamic transients associated with expansion. Bubbles are identified through handscanning of images, counting bubbles that are nucleated in the visible volume and are coincident with the bubble chamber trigger. No excess rate of bubble nucleation at the wall of the chamber was observed, so no fiducial cut is taken.
@@ III.4 Coincident Scintillation Veto
While the scintillation signal recorded by the PMT in principle could allow for energy reconstruction of NR events, in practice the photon yield for the (keV) events of interest here is too low to be useful, giving an average of photon detected in XeBC. However, the PMT signal still provides a useful method of background suppression, easily tagging alpha-decays and transiting charged particles from cosmic-ray showers, which dominate the background bubble nucleation rate.
The scintillation trigger is set to be sensitive to single photoelectrons in the PMT with % efficiency in all data where the scintillation veto is used, verified using a pulsed blue LED. Light collection efficiency is measured by exposing the chamber to 122-keV -rays from a 57Co source. The 122 keV photoabsorption peak is detected at 4 photoelectrons, corresponding to a total photon detection efficiency of 0.05%. This efficiency is approximately a factor of 8 lower than reported in Ref. [12], and is thought to be due to decreased UV transparency in the new quartz vial. With this efficiency, nuclear recoils around 120 keV result in 1 photoelectron collected on average, based on the zero-field nuclear recoil light yields given in NEST v2.2.0 [34].
@@
Figure 2 shows the scintillation spectra associated with single- and multi-bubble events in 252Cf and background data. The 252Cf source emits neutrons with energies up to 10 MeV [29], giving an elastic NR energy spectrum that peaks well below the 120 keVnr light-collection threshold. Larger scintillation pulses of tens of photoelectrons accompany inelastic scatters, where the scintillation is dominated by ERs produced through nuclear relaxation. Background single-bubble events from -decays and cosmic rays give much stronger scintillation, with 85% of background events coincident with scintillation pulses of photoelectrons. A cut at 512 photoelectrons reduces the single-bubble background rate from 0.2 bubbles/minute to 0.02 bubbles/minute without impacting nuclear recoils from 252Cf.

### v2

@@ Abstract
A device filled with pure xenon first demonstrated the ability to operate simultaneously as a bubble chamber and scintillation detector in 2017. Initial results from data taken at thermodynamic thresholds down to 4 keV showed sensitivity to 20 keV nuclear recoils with no observable bubble nucleation by -ray interactions. This paper presents results from further operation of the same device at thermodynamic thresholds as low as keV, hardware limited. The bubble chamber has now been shown to have sensitivity to 1 keV nuclear recoils while remaining insensitive to bubble nucleation by -rays. A data-driven calibration of the chamber’s nuclear recoil nucleation response, as a function of nuclear recoil energy and thermodynamic state, is presented. Stringent upper limits are established for the probability of bubble nucleation by -ray-induced Auger cascades, with a limit of set at keV, the lowest thermodynamic threshold explored.
@@ I Introduction
It was shown in Ref. [12] that pure xenon bubble chambers are sensitive to nuclear recoils, as necessary for a dark matter search, despite being insensitive to -rays. Thus, the scintillating noble liquid bubble chamber is a promising technology for exploring dark matter, in particular (GeV) dark matter where sensitivity to sub-keV nuclear recoils is necessary, as well as for detection of coherent elastic neutrino-nucleus scattering (CENS) [13, 14]. This paper describes further experiments with the same xenon bubble chamber (referred to as “XeBC”) as the one used in Ref. [12], modified for operation at higher degrees of superheat (lower nuclear recoil detection thresholds). A variety of -ray and neutron sources are used to characterize the response of superheated liquid xenon to both electronic and nuclear recoils. Preliminary results from the data presented here [15, 16] have already motivated the construction of two liquid argon scintillating bubble chambers currently being installed at Fermilab and SNOLAB by the SBC Collaboration [17, 18]. The calibration strategies and analysis techniques presented in this work are directly relevant to these chambers, and to other future dark matter direct detection and CENS experiments that rely on measurements with neutron and -ray sources to constrain the detector’s low-energy response.
@@ II Previous Experiments and Motivation
The PICO Collaboration has recently published extensive studies of electronic recoil (ER) and nuclear recoil (NR) induced bubble nucleation in C3F8 bubble chambers [19, 20, 21, 22]. The well-established “Seitz” threshold [3, 19] (calculated from the pressure, temperature, and thermophysical properties of the target fluid, and denoted here as ) is used both as a measure of the degree of superheat in these studies, and as an indication of the expected NR detection threshold. In practice, this quantity is seen to underestimate both observed [22, 23, 24, 25, 26, 27] and simulated [28, 29] thresholds for NR-induced bubble nucleation. The level of disagreement varies with both target fluid and threshold, requiring NR nucleation efficiencies to be determined in each new case through dedicated neutron calibrations. The data-driven approach developed by the PICO Collaboration [22] consists of a global fit of simulated to observed bubble nucleation rates with data gathered from multiple neutron calibration sources.
@@
In 2017, XeBC observed for the first time coincident bubble nucleation and scintillation from NRs [12] without the bubble nucleation by photoabsorption seen in xenon-contaminated C3F8. This paper presents new results from low-threshold (high-superheat) XeBC data taken in the year following the 2017 publication, as described in Sec. III, with NR and ER analyses in the style of the above PICO analyses presented in Sec. IV. Section IV.1 presents constraints on the NR nucleation efficiency in xenon at Seitz thresholds between 0.9 and 2.1 keV. This analysis takes an approach similar to Ref. [22], making minimal assumptions regarding the dependence of nucleation efficiency on either recoil energy or thermodynamic state. Section IV.2 gives upper limits on XeBC’s sensitivity to ER events at Seitz thresholds down to 0.5 keV. No evidence for bubble nucleation by -ray photoabsorption is seen, extending the disparity between xenon and the molecular fluids studied by PICO.
@@ III.2 Chamber control and instrumentation
The thermodynamic state (temperature and pressure) of a bubble chamber determines the threshold energy required for bubble nucleation. For XeBC, there are two temperature regions to consider: the warm active region becomes superheated when the chamber pressure drops to the target expanded state, while the jar seal, plumbing, and the bellows that transmits pressure changes from the hydraulic fluid to the target fluid are kept cold so that liquid contacting these surfaces will not boil [20]. The entire chamber assembly is cooled by a cold finger running from an open-mouth dewar of liquid nitrogen to the top of the cold (inactive) region of the bubble chamber. The warm region is then heated with a nichrome wire heater mounted on the outer wall of the copper, aluminum, and sapphire radiation shield that encloses the target volume. This shield is thermally coupled to the outer jar by an indium contact immediately below the target volume. The temperature of the warm region is controlled with feedback from a 100- platinum RTD mounted on the inside of the radiation shield (T2 in Fig. 1). Additional RTDs (T1, T3, T4) monitor temperatures throughout the warm and cold regions.
@@
Pressure is controlled by a hydraulic system filled with Dynalene-MV, a commercial low-temperature / low-viscosity hydrocarbon blend, using a pair of hydraulic accumulators as “reservoirs” for the high (compressed) and low (expanded) pressures. Hydraulic pressure is transmitted to the xenon by an in-line pressure-balanced bellows. A second bellows transmits the pressure back to a warm hydraulic stub, and the pressure of all three volumes (hydraulic system output, xenon, and hydraulic stub) are monitored by absolute pressure transducers with 0.01 bar resolution (PT5, PT4, and PT6, respectively). The redundant pressure readings provide cross-calibrations to compensate for temperature-dependent readings in the cold xenon pressure transducer and for pressure differentials across the bellows due to the bellows spring constants.
@@
Bubble chamber data is organized by “event,” i.e., one cycle from compressed (insensitive) expanded (superheated, sensitive) trigger compressed. During a run the chamber is set to cycle automatically, repeating the cycle with a 60 s compression period between events. The camera acts as the primary bubble trigger, collecting 0.3-megapixel, 8-bit images at 100 frames per second with synchronized illumination from a ring of 950 nm LEDs mounted around the camera. The frame grabber receiving the camera images broadcasts a trigger when a significant change (typ. 25% full scale) is seen in a minimum number of pixels (typ. 12 pixels) between consecutive images. In addition to images, data recorded for each event include pressure and temperature readings throughout the event, acoustic waveforms digitized at 2.5 MHz covering 100 ms before and after the bubble trigger, and time-stamped PMT waveforms digitized at 1 GHz for scintillation pulses observed throughout the event. The PMT digitizer is triggered by a Nuclear Instrumentation Module (NIM) discriminator. This trigger is suppressed when camera illumination is on, incurring a 10% dead time in the scintillation channel, but is otherwise independent from the bubble chamber cycle.
@@ III.3 Low-threshold data
The low-threshold data analyzed here were acquired over the course of five months from June to October 2017, in three distinct xenon fills. The chamber’s superheated operational temperature ranged from to , with target expansion pressures from 1.7–3.4 bara. Data are reported in terms of the Seitz threshold of the superheated target, calculated following [19] with thermophysical properties for xenon taken from NIST REFPROP [31]. Pressure and temperature were controlled in this dataset to a precision of at least 0.04 bar and 0.1∘C, corresponding to uncorrelated uncertainties on of 0.8% and 1.2%, respectively, across datasets. Because xenon temperature was measured only indirectly (by measuring the temperature of external elements thermally coupled to the target volume), there is an additional 1∘C global temperature uncertainty, with a corresponding 12% global uncertainty on . The uncorrelated uncertainties on are small compared to statistical uncertainties in the data presented below, while the only effect of the global uncertainty on is an effective relabeling of fluid states. Conclusions regarding actual nuclear recoil detection thresholds and electron recoil sensitivities are unaffected, except for the conservative constraint imposed in the analysis that there is zero sensitivity to nuclear recoils with energies below . For these reasons, uncertainties on the thermal state of the chamber are neglected in the analysis.
A variety of sources were deployed, including 252Cf as a fission neutron source, 88Y-Be and 207Bi-Be as photoneutron sources, and 88Y, 207Bi, and 57Co as -ray sources. Table 1 summarizes the activities, neutron production rates, placement, and resulting neutron fluxes in the superheated target from these sources as determined by an MCNPX-PoliMi [32] Monte Carlo simulation. The fission source generates neutrons over a broad spectrum with an average energy of 2.3 MeV [33], and is simulated using the built-in PoliMi 252Cf fission spectrum. The photoneutron sources produce nearly monoenergetic low-energy neutrons via 9BeBe reactions in BeO and metallic beryllium parts surrounding the -ray sources. The 88Y-Be source neutrons are peaked at 152 keV, with approximately 0.7% of the neutrons at higher energy (950 keV) from the low-branching-fraction 2734-keV -ray [34]. The 207Bi-Be source produces neutrons peaked at 91 keV. Photoneutron rates and neutron spectra are calculated using MCNPX-PoliMi, using the 9BeBe cross sections given in [35] and correcting lab-frame neutron energies with the appropriate relativistic kinematic factor (necessary due to MCNPX’s mishandling of momentum [36]).
@@
Table 1 also shows uncertainties in neutron flux from each source, stemming from uncertainties both in source strength and in neutron propagation. The 252Cf source neutron emission rate was calibrated to 6% precision with a well-characterized 3He counter, and was found to be consistent with the high-energy (2 MeV) gamma emission rate measured by a NaI crystal. The total uncertainty on neutron flux from the 252Cf source at the chamber is 9% after including geometric uncertainties in neutron propagation, which were estimated by varying selected details in the simulated geometry and by comparing the MCNP-PoliMi simulation with an independently developed Geant4 [37] simulation. The uncertainty in the 88Y-Be source neutron emission rate is 5%, due to a combination of geometrical uncertainties in the 88Y deposition profile in the source capsule and uncertainties on the 9BeBe cross section, and is consistent with the neutron emission rate measured by the same 3He counter used to characterize the 252Cf source. Neutron propagation uncertainties are significantly higher for 88Y-Be than for 252Cf, giving a total neutron flux uncertainty of 13%. The 207Bi-Be neutron flux uncertainty is larger still at 30%, dominated by uncertainty associated with additional beryllium pieces deployed with that source in an attempt to boost the neutron emission rate. Finally, there is a common-mode 10% uncertainty on the integrated flux of neutrons through the “active” xenon due to uncertainty on the volume of xenon visible to the camera.
@@
Each neutron dataset (characterized by a single source and threshold) is accompanied by a dedicated background measurement taken at the same threshold and in the same xenon fill. In the case of 252Cf this simply involves removing the source. For the 88Y-Be and 207Bi-Be experiments, the Be and BeO components are replaced with aluminum, with the 88Y and 207Bi sources still present to preserve the same -ray flux incident upon the detector but without the photoneutrons. The 207Bi and 88Y -ray sources without beryllium are also used to measure ER sensitivity, along with the lower-energy 57Co source, at Seitz thresholds between 0.50 keV and 1.56 keV. Sources were placed 10 cm from the active xenon for all ER sensitivity measurements.
@@
The reduced data for the subsequent analysis includes only live times and bubble counts at each threshold and source configuration (with the exception of the 252Cf data, which includes the scintillation veto described in the next section). The chamber is considered live only when the measured pressure is within 0.034 bar (0.5 psi) of the target pressure set-point (extended to 0.068 bar for the 0.90 keV 88Y-Be dataset), and any live time in the the first 20 s of each event is discarded to avoid thermodynamic transients associated with expansion. Bubbles are identified through hand-scanning of images, counting bubbles that are nucleated in the visible volume and are coincident with the bubble chamber trigger. No excess rate of bubble nucleation at the wall of the chamber was observed, so no fiducial cut is taken.
@@ III.4 Coincident scintillation veto
While the scintillation signals produced by NR events can in principle be used for energy reconstruction, in practice NR photon yields are too low to be useful for this purpose. Both in XeBC and in SBC’s planned dark matter and CENS detectors, any practical light collection configuration yields photon detected for the (keV) and sub-keV nuclear recoils in these experiments’ regions of interest. However, the scintillation signal still provides a useful method of background suppression, easily tagging alpha decays and transiting charged particles from cosmic-ray showers, which dominate the background bubble nucleation rate in XeBC.
The scintillation trigger is set to be sensitive to single photoelectrons in XeBC’s single PMT with % efficiency in all data where the scintillation veto is used, verified using a pulsed blue LED. Light collection efficiency is measured by exposing the chamber to 122-keV -rays from a 57Co source. The 122 keV photoabsorption peak is detected at 4 photoelectrons, corresponding to a total photon detection efficiency of 0.05%. This efficiency is approximately a factor of 8 lower than reported in Ref. [12], and is thought to be due to decreased UV transparency in the new quartz vial. With this efficiency, nuclear recoils around 120 keV result in one photoelectron collected on average, based on the zero-field nuclear recoil light yields given in NEST v2.2.0 [38].
@@
Figure 2 shows the scintillation spectra associated with single- and multiple-bubble events in 252Cf and background data. The 252Cf source emits neutrons with energies up to 10 MeV [33], giving an elastic NR energy spectrum that peaks well below the 120 keVnr light collection threshold. Larger scintillation pulses of tens of photoelectrons accompany inelastic scatters, where the scintillation is dominated by ERs produced through nuclear relaxation. Background single-bubble events from -decays and cosmic rays give much stronger scintillation, with 85% of background events coincident with scintillation pulses of photoelectrons. A cut at 512 photoelectrons reduces the single-bubble background rate from 0.2 bubbles/minute to 0.02 bubbles/minute without impacting nuclear recoils from 252Cf.